A farmer has enough food to feed 28 animals in his cattle for 9 days. How long will the food last, if there were 8 more animals in his cattle?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
In this particular question use the property that if there are x animals and food last for n days and if there are y number of animals increased then the number of days the food will last is the ratio of the product of the x animals and n number of days to the total number of animals (i.e. x + y) so use this concept to reach the solution of the question. Complete step-by-step answer: Given data: There are 28 animals a farmer has. Now the farmer has enough food to feed the 28 animals for 9 days. Now if the number of animals increased by 8 then we have to calculate the number of days the food lasts. So the total number of animals is the sum of the previous animals and the increased animals. So the total number of animals is = 28 + 8 = 36 animals. Now the number of days the food will last is the ratio of the multiplication of the previous number of animals and the previous number of days the food will last to the total number of animals. So the number of days the food will last is = 28×936 Now simplify the above equation we have, So the number of days the food will last is = 284=7 days. So the food will last now for 7 days. So this is the required answer. Note – Whenever we face such types of questions the key concept we have to remember is that always recall the formula of the number of days the food will last if the number of animals is increased which is stated above so use this formula simply substitute the values as above and simplify we will get the required answer.